# On Geometric Evolution and Microlocal Regularity of the Navier–Stokes Equations

### v1

@@ Abstract
We develop a geometric and microlocal framework for the Navier–Stokes equations by lifting the dynamics to the cosphere bundle of a Riemannian manifold. In this formulation, the velocity field and vorticity are represented as microlocal distributions whose evolution is governed by a transport–dissipation system generated by a canonical dynamical vector field described with differential forms. We introduce microlocal amplitudes, directional energy functionals, and monotone volume invariants on the compact phase space, which quantify directional concentration and alignment mechanisms associated with potential loss of regularity. The viscous term induces effective geometric diffusion on the cosphere bundle, yielding closed differential inequalities in a geometric setting. To capture the interaction between fluid deformation and geometry, we define an effective connection and curvature tensor encoding the influence of the symmetric velocity gradient. This structure gives rise to a Ricci-type geometric evolution at the microlocal level, controlled by a dissipative functional that excludes extreme directional concentration and persistent transverse stretching. Our results do not resolve the global regularity problem in full generality, but provide a coherent geometric mechanism that severely restricts blow-up scenarios compatible with viscous Navier–Stokes dynamics, reformulating the analytic regularity question as a problem of dissipative stability on a compact phase space.
@@ III Microlocal Lifting and the Cosphere Bundle
where $\alpha$ is a fixed geometric coupling parameter [21]. This metric deformation encodes the interaction between the background geometry and the fluid deformation tensor.
@@
$\quad-(\Gamma^{m}_{ik}+g^{\prime}_{ik}u^{m})(\Gamma^{l}_{jm}+g^{\prime}_{jm}u^{l}).$ (36)
@@
The effective Ricci tensor is obtained by contraction,
@@
$R^{u}_{jk}\simeq R_{jk}+\alpha\,\nabla^{i}\bigl(\nabla_{j}S_{ik}+\nabla_{k}S_{ij}-\nabla_{i}S_{jk}\bigr).$ (42)
@@ IV Geometric Evolution Operators and Microlocal Representation of Solutions
where $\chi$ is a smooth cutoff function. Restricting by homogeneity to $S^{*}M$, we obtain $\tilde{\phi}\in C^{\infty}(S^{*}M)$.
@@
$\partial_{t}\tilde{\phi}+\mathcal{L}_{X_{u}}\tilde{\phi}+\nu\|\xi\|_{g}^{2}\,\tilde{\phi}=\tilde{R}+\tilde{F},\qquad\langle\tilde{\phi},\xi\rangle_{g}=0$ (52)
@@
$\Pi_{\xi}(\eta)=\eta-\frac{\langle\eta,\xi\rangle_{g}}{\|\xi\|_{g}^{2}}\xi$ (54)
@@
This causes the pressure term to vanish when restricting the equation to the cosphere bundle and projecting onto transverse directions. In fact, the condition of the divergence transform also justifies the defined projection and confirms that the dynamics evolve naturally in $S^{*}M$.
In order to solve the microlocal system on the cosphere bundle $S^{*}M$, we define the Green operator $G(t,s)$ as the fundamental solution of the non-autonomous dissipative equation [25, 38]. Let $Q(t)=\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2}$ be the differential operator acting on sections of the bundle. The Green operator satisfies
$\left(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2}\right)G(t,s)=\delta(t-s)\mathbb{I}$ (56)
@@
$G(t,s)[\tilde{f}](x,\xi)=\exp\left(-\int_{s}^{t}\nu\|\xi(\tau)\|_{g}^{2}\,d\tau\right)\tilde{f}(\Psi_{s,t}(x,\xi))$ (57)
@@
$(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2})G(t,s)\\ =(-\nu\|\xi\|_{g}^{2}+\nu\|\xi\|_{g}^{2}+\mathcal{L}_{X_{u}}-\mathcal{L}_{X_{u}})G(t,s)=0$ (59)
@@
$\tilde{\phi}(t)=\mathcal{T}\bigg[\exp\left(-\int_{t_{0}}^{t}\big(\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2}\big)\,ds\right)\tilde{\phi}_{0}\bigg]\\ +\int_{t_{0}}^{t}G(t,s)\big[\tilde{R}(s)+\tilde{F}(s)\big]\,ds$ (60)
@@
which defines a geometrically canonical reconstruction of the velocity field.
@@
$\partial_{t}\tilde{\omega}+\mathcal{L}_{X_{u}}\tilde{\omega}=-\nu\|\xi\|_{g}^{2}\tilde{\omega}+\tilde{R}_{\omega}+i\xi\wedge\tilde{F},\qquad\xi\wedge\tilde{\omega}=0$ (62)
@@
$\tilde{\omega}(t)=\mathcal{T}\bigg[\exp\left[-\int_{t_{0}}^{t}(\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2})\,ds\right]\tilde{\omega}_{0}\bigg]\\ +\int_{t_{0}}^{t}G(t,s)\big[\tilde{R}\omega(s)+i\xi\wedge\tilde{F}(s)\big]ds$ (63)
If $\tilde{\phi}(t)$ is a solution to (60), then $\tilde{\omega}(t):=d\tilde{\phi}(t)$ satisfies the microlocal vorticity equation (62). This is because the exterior derivative commutes with the differential operators used in the system for smooth differential forms, and $d(dP)=0$; therefore, applying the exterior derivative to (52) yields exactly (62) with a solution reconstructed also from
@@
It is inferred that the microlocal lift of the Navier-Stokes equations produces a well-defined dissipative dynamical system on the cosphere bundle, whose solution admits an explicit representation through evolution operators. The exact compatibility between the velocity and vorticity equations guaranties the consistency of the formalism. This approach reduces the global regularity problem to the control and convergence of geometric microlocal invariants, the analysis of which is addressed in the following section.
@@ V Stability, Energy and Geometric Criteria for Regularity and Smoothness
Throughout this section, $\lambda\in\{\phi,\omega\}$ denotes either the velocity $1$-form or the vorticity $2$-form on $M$, and $\tilde{\lambda}$ its microlocal lift to $S^{*}M$. All norms and contractions are taken with respect to the effective metric $g^{\prime}$ introduced earlier. Since $S^{*}M$ is compact and endowed with a canonical contact structure, all integrals below are well-defined without further assumptions.
We begin by defining the fundamental local geometric quantity governing the analysis. The microlocal amplitude of $\tilde{\lambda}$ is defined pointwise by the contraction
$|\tilde{\lambda}|_{g^{\prime}}^{2}:=\langle\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}=\iota_{\tilde{\lambda}^{\sharp}}\tilde{\lambda}$, (65)
where $\tilde{\lambda}^{\sharp}$ denotes the metric dual with respect to $g^{\prime}$. This quantity measures the intensity of the lifted differential form along a given cotangent direction. Unlike classical norms, it retains full directional information encoded in the cosphere bundle.
The corresponding global geometric functional is obtained by pairing this density with the canonical Liouville volume form on $S^{*}M$,
@@
where $\alpha$ is the contact form induced from the Liouville $1$-form on $T^{*}M$. We define the microlocal energy functional
$\mathcal{V}_{\lambda}(t):=\int_{S^{*}M}|\tilde{\lambda}(x,\xi,t)|_{g^{\prime}}\,\mu_{S^{*}M}$ (67)
Geometrically, $\mathcal{V}_{\lambda}$ measures the total directional spread of the field $\lambda$ in phase space. Physically, it quantifies how much the solution concentrates along preferred cotangent directions, independently of its spatial localization.
We now derive a balance law for $\mathcal{V}_{\lambda}$. The microlocal evolution equation satisfied by $\tilde{\lambda}$ is
$\partial_{t}\tilde{\lambda}+\mathcal{L}_{X_{u}}\tilde{\lambda}+\nu\|\xi\|_{g}^{2}\,\tilde{\lambda}=\tilde{R}[\tilde{\lambda}]+\tilde{C}[\tilde{\lambda}]+\tilde{F}$ (68)
where $X_{u}$ is the canonical lifted transport field on $S^{*}M$. We take the pointwise inner product with $\tilde{\lambda}$ and use bilinearity to obtain
$\frac{1}{2}\partial_{t}|\tilde{\lambda}|_{g^{\prime}}^{2}+\frac{1}{2}\mathcal{L}_{X_{u}}|\tilde{\lambda}|_{g^{\prime}}^{2}+\nu\|\xi\|_{g}^{2}|\tilde{\lambda}|_{g^{\prime}}^{2}\\ =\langle\tilde{R}[\tilde{\lambda}],\tilde{\lambda}\rangle+\langle\tilde{C}[\tilde{\lambda}],\tilde{\lambda}\rangle+\langle\tilde{F},\tilde{\lambda}\rangle$ (69)
Each term on the right-hand side admits a sharp geometric bound. Since the curvature term is zeroth order in $\xi$,
$|\langle\tilde{R}[\tilde{\lambda}],\tilde{\lambda}\rangle|\leq\|\mathrm{Ric}\|_{L^{\infty}(M)}\,\|\tilde{\lambda}\|_{g^{\prime}}^{2}$ (70)

### v2

@@ Abstract
We develop a geometric and microlocal framework for the Navier–Stokes equations by lifting the dynamics to the cosphere bundle of a Riemannian manifold. In this formulation, the velocity field and vorticity are represented as microlocal distributions whose evolution is governed by a linear transport–dissipation system generated by a canonical dynamical vector field. We introduce microlocal amplitudes, directional energy functionals, and monotone volume invariants on the compact phase space, which quantify directional concentration and alignment mechanisms associated with potential loss of regularity. The viscous term induces an effective geometric diffusion on the cosphere bundle, yielding closed differential inequalities in a geometric setting. To capture the interaction between fluid deformation and geometry, we define an effective connection and curvature tensor encoding the influence of the symmetric velocity gradient. This structure gives rise to a Ricci–type microlocal geometric evolution that constrains directional stretching and excludes extreme angular concentration compatible with viscous dissipation. While the present results do not resolve the global regularity problem, they provide a coherent geometric mechanism that severely restricts admissible blow–up scenarios, reformulating the regularity question as a problem of dissipative stability on a compact phase space.
@@ I Introduction: The Regularity Problem and Geometric Flows
Microlocal lifting approach
The microlocal lifting to the cosphere bundle $S^{*}M$ adopted in this paper should be understood as a parametric refinement and compactification of the dynamics on the base manifold $M$. All geometric objects—including the effective connection $\nabla^{u}$, its torsion and curvature, and the metric $g^{\prime}$—are defined on $M$, with coefficients depending smoothly on $(x,\xi)\in S^{*}M$. The cotangent variable $\xi$ serves to encode directional and spectral information, but no geometric structure is imposed on the tangent bundle of $S^{*}M$ itself. This viewpoint allows us to combine microlocal analysis with intrinsic geometric evolution without introducing additional bundle-level connections.
@@ III Microlocal Lifting and the Cosphere Bundle
where $\alpha$ is a fixed geometric coupling parameter [21]. This metric deformation encodes the interaction between the background geometry and the fluid deformation tensor. We define the effective covariant derivative $\nabla^{u}$ acting on vector fields by
$(\nabla^{u})_{i}v^{k}:=\nabla_{i}v^{k}+A^{k}{}_{ij}v^{j}$ (36)
where $\nabla$ denotes the Levi–Civita connection of $g$ and $A^{k}{}_{ij}$ is a (1,2)(1,2)–tensor encoding the deformation induced by $u$.
The effective connection $\nabla^{u}$ is, in general, not metric–compatible with respect to the effective metric $g^{\prime}$. Metric compatibility would require
$(\nabla^{u}_{i}g^{\prime})_{jk}=0$ (37)
where, by definition,
$(\nabla^{u}_{i}g^{\prime})_{jk}:=\partial_{i}g^{\prime}_{jk}-(\Gamma^{u})^{\ell}{}_{ij}\,g^{\prime}_{\ell k}-(\Gamma^{u})^{\ell}{}_{ik}\,g^{\prime}_{j\ell}$ (38)
Using the decomposition $(\Gamma^{u})^{\ell}{}_{ij}=\Gamma^{\ell}{}_{ij}+A^{\ell}{}_{ij}$ and recalling that $\nabla$ denotes any connection, for example, using the Levi–Civita connection associated with the background metric $g$, we obtain
$(\nabla^{u}_{i}g^{\prime})_{jk}$ $=(\nabla_{i}g^{\prime})_{jk}-A^{\ell}{}_{ij}\,g^{\prime}_{\ell k}-A^{\ell}{}_{ik}\,g^{\prime}_{j\ell}$ (39)
In particular, if $g^{\prime}$ is covariantly constant with respect to $\nabla$, that is $\nabla g^{\prime}=0$, the expression reduces to
$(\nabla^{u}_{i}g^{\prime})_{jk}=-A^{\ell}{}_{ij}\,g^{\prime}_{\ell k}-A^{\ell}{}_{ik}\,g^{\prime}_{j\ell}$ (40)
Equivalently, in invariant tensorial form,
$\nabla^{u}g^{\prime}=-\bigl(A(u)\cdot g^{\prime}+g^{\prime}\cdot A(u)\bigr)$ (41)
where $A(u)$ is viewed as a (1,1)(1,1)–tensor acting on covariant indices.
Therefore, the effective connection $\nabla^{u}$ is metric–compatible with respect to $g^{\prime}$ if and only if the deformation tensor $A(u)$ is skew–adjoint with respect to $g^{\prime}$, namely if
$g^{\prime}[A(u)X,Y]+g^{\prime}[X,A(u)Y]=0$ (42)
Here $X,Y$ denote smooth vector fields on the base manifold $M$, such that, $X,Y\in\Gamma(TM)$. Although the tensors $A(u)$ and $g^{\prime}$ may depend microlocally on $(x,\xi)\in S^{*}M$, the effective connection $\nabla^{u}$ acts on tensor fields over $M$, with the cotangent variables playing the role of parameters.
In general, this condition is violated by the symmetric part of $\nabla u$, and $\nabla^{u}$ defines a non–metric affine connection encoding the velocity–induced deformation of the effective geometry encoding the transport induced by the velocity field controlled by $\|\nabla u\|_{L^{\infty}}$, which is sufficient for the energy estimates derived below. Then, the canonical decomposition considered earlier remains orthogonal with respect to the background metric $g$ and orthogonal with respect to the effective metric $g^{\prime}$ up to lower-order corrections.
Since the Levi–Civita connection is torsion–free, the torsion tensor $T^{u}$ of $\nabla^{u}$ is entirely determined by the antisymmetric part of $A$. Indeed, by definition,
$(T^{u})^{k}{}_{ij}$ $:=(\Gamma^{u})^{k}{}_{ij}-(\Gamma^{u})^{k}{}_{ji}$ (43)
[…]
Hence,
$(T^{u})^{k}{}_{ij}=A^{k}{}_{ij}-A^{k}{}_{ji}$ (46)
We now specify the structure of the deformation tensor $A$. We decompose it as
$A^{k}{}_{ij}=g^{\prime k\ell}g^{\prime}_{im}\nabla_{j}u^{m}+B^{k}{}_{ij}$ (47)
where $g^{\prime}$ is the effective metric introduced previously and $B^{k}{}_{ij}$ is symmetric in the lower indices,
$B^{k}{}_{ij}=B^{k}{}_{ji}$ (48)
Substituting this decomposition into the expression for the torsion, the symmetric part cancels and we obtain
[…]
$=g^{\prime k\ell}\left(g^{\prime}_{im}\nabla_{j}u^{m}-g^{\prime}_{jm}\nabla_{i}u^{m}\right)$ (50)
This computation shows that
  • the torsion $T^{u}$ is antisymmetric in the indices $i,j$.
  • $T^{u}$ is of first order in $\nabla u$.
  • $T^{u}\equiv 0$ if and only if the velocity field $u$ is covariantly constant.
In particular, the effective connection $\nabla^{u}$ is torsion–free if and only if the velocity field $u$ is covariantly constant.
@@
$\quad-(\Gamma^{m}_{ik}+g^{\prime}_{ik}u^{m})(\Gamma^{l}_{jm}+g^{\prime}_{jm}u^{l})$ (51)
@@
The curvature tensor $(R^{u})^{l}{}_{ijk}$ associated with the effective connection $\nabla^{u}$ satisfies a modified second Bianchi identity, reflecting the presence of torsion-like terms induced by the velocity-dependent deformation of the connection. More precisely, since the effective connection $\nabla^{u}$ is an affine connection with torsion, its curvature tensor satisfies the standard second Bianchi identity for connections with torsion. In index notation, one has
$\nabla^{u}_{[m}(R^{u})^{l}{}_{|i|jk]}=(T^{u})^{p}{}_{[mj}(R^{u})^{l}{}_{|i|k]p}$ (55)
where $(T^{u})^{p}{}_{ij}$ denotes the torsion tensor of $\nabla^{u}$, defined earlier. Here, brackets $\cdot$] denote antisymmetrization over the indicated indices, while vertical bars $|\cdot|$ exclude the enclosed index from the antisymmetrization. This identity follows directly from the definition of the curvature operator $\forall\ X,Y,Z\in\Gamma(TM)$, satisfying,
$R^{u}(X,Y)Z=\nabla^{u}_{X}\nabla^{u}_{Y}Z-\nabla^{u}_{Y}\nabla^{u}_{X}Z-\nabla^{u}_{[X,Y]}Z$ (56)
together with the Jacobi identity for iterated covariant derivatives [30]. The additional torsion term on the right-hand side of (55) measures precisely the failure of $\nabla^{u}$ to be torsion-free, and vanishes identically when $T^{u}\equiv 0$.
Since the deformation tensor $A$ depends smoothly on the velocity field $u$ and its first derivatives, the torsion tensor $(T^{u})^{p}{}_{ij}$ is of first order in $\nabla u$. Consequently, the right-hand side of (55) consists of lower-order terms which are uniformly controlled under the regularity assumptions imposed on $u$. In particular, the classical second Bianchi identity is recovered in the limit where $A(u)$ vanishes. In the general case, the deviation from the classical identity is entirely encoded in controlled lower-order terms involving $\nabla u$.
Now, the effective Ricci tensor is obtained by contraction,
@@
$R^{u}_{jk}\simeq R_{jk}+\alpha\,\nabla^{i}\bigl(\nabla_{j}S_{ik}+\nabla_{k}S_{ij}-\nabla_{i}S_{jk}\bigr)$ (59)
@@ IV Geometric Evolution Operators and Microlocal Representation of Solutions
where $\chi$ is a smooth cutoff function. Restricting by homogeneity to $S^{*}M$, we obtain $\tilde{\phi}\in C^{\infty}(S^{*}M)$. The transform (68) should be understood as a directional microlocal projection rather than a pointwise invertible transform, because of that, the horizontal lift is uniquely determined once the effective connection is fixed, although it is not canonical with respect to the background metric.
@@
$\partial_{t}\tilde{\phi}+\mathcal{L}_{X_{u}}\tilde{\phi}+\nu\|\xi\|_{g^{\prime}}^{2}\,\tilde{\phi}=\tilde{R}+\tilde{F},\qquad\langle\tilde{\phi},\xi\rangle_{g^{\prime}}=0$ (69)
@@
$\Pi_{\xi}(\eta)=\eta-\frac{\langle\eta,\xi\rangle_{g^{\prime}}}{\|\xi\|_{g^{\prime}}^{2}}\xi$ (71)
@@
This causes the pressure term to vanish when restricting the equation to the cosphere bundle and projecting onto transverse directions because it acts on the base manifold variables and does not generate transverse components in phase space. In fact, the condition of the divergence transform also justifies the defined projection and confirms that the dynamics evolve naturally in $S^{*}M$.
In order to solve the microlocal system on the cosphere bundle $S^{*}M$, we define the Green operator $G(t,s)$ as the fundamental solution of the non-autonomous dissipative equation [25, 38]. Let $Q(t)=\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}$ be the differential operator acting on sections of the bundle. The Green operator satisfies
$\left(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}\right)G(t,s)=\delta(t-s)\mathbb{I}$ (73)
@@
$G(t,s)[\tilde{f}](x,\xi)=\exp\left(-\int_{s}^{t}\nu\|\xi(\tau)\|_{g^{\prime}}^{2}\,d\tau\right)\tilde{f}(\Psi_{s,t}(x,\xi))$ (74)
@@
$(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2})G(t,s)\\ =(-\nu\|\xi\|_{g^{\prime}}^{2}+\nu\|\xi\|_{g^{\prime}}^{2}+\mathcal{L}_{X_{u}}-\mathcal{L}_{X_{u}})G(t,s)=0$ (76)
@@
$\tilde{\phi}(t)=\mathcal{T}\bigg[\exp\left(-\int_{t_{0}}^{t}\big(\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}\big)\,ds\right)\tilde{\phi}_{0}\bigg]\\ +\int_{t_{0}}^{t}G(t,s)\big[\tilde{R}(s)+\tilde{F}(s)\big]\,ds$ (77)
@@
which defines a geometrically canonical reconstruction of the velocity field. The reconstruction equation (78) recovers $\phi$ in a weak or averaged sense, which is sufficient for the purposes of the present analysis.
@@
$\partial_{t}\tilde{\omega}+\mathcal{L}_{X_{u}}\tilde{\omega}=-\nu\|\xi\|_{g^{\prime}}^{2}\tilde{\omega}+\tilde{R}_{\omega}+i\xi\wedge\tilde{F},\qquad\xi\wedge\tilde{\omega}=0$ (79)
@@
$\tilde{\omega}(t)=\mathcal{T}\bigg[\exp\left[-\int_{t_{0}}^{t}(\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2})\,ds\right]\tilde{\omega}_{0}\bigg]\\ +\int_{t_{0}}^{t}G(t,s)\big[\tilde{R}\omega(s)+i\xi\wedge\tilde{F}(s)\big]ds$ (80)
If $\tilde{\phi}(t)$ is a solution to (77), then $\tilde{\omega}(t):=d\tilde{\phi}(t)$ satisfies the microlocal vorticity equation (79). This is because $d(dP)=0$ and $X_{u}$ depends on $u$ and its derivatives, the exterior derivative does not commute exactly with the Lie derivative along $X_{u}$, a direct computation yields the commutator identity
$d,\mathcal{L}_{X_{u}}]=\mathcal{L}_{dX_{u}}$ (81)
When acting on microlocal scalar amplitudes $\tilde{\phi}$, the right–hand side produces lower–order terms depending on $\nabla u$, and we obtain the estimate
$d,\mathcal{L}_{X_{u}}]\tilde{\phi}=\mathcal{O}(\nabla u)\tilde{\phi}$ (82)
The viscous contribution in the lifted equation is modeled by the multiplication operator
$D:=\nu\|\xi\|_{g^{\prime}}^{2}$ (83)
whose commutator with $d$ is given explicitly by
$d,D]=d(\nu\|\xi\|_{g^{\prime}}^{2})\wedge(\cdot)$ (84)
Combining the transport and dissipative contributions, we define the total commutator operator
$\left[d,\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}\right]={R}$ (85)
which is a zeroth–order operator controlled by the geometry and the velocity field. More precisely, there exists a constant $C>0$ such that
$\|{R}_{u}\|_{L^{2}\to L^{2}}\leq C\left(\|\nabla u\|_{L^{\infty}}+1\right)$ (86)
As a consequence, if $\tilde{\phi}$ is a solution of the lifted scalar equation, then the microlocal vorticity $\tilde{\omega}:=d\tilde{\phi}$ satisfies the corresponding lifted equation up to lower–order controlled terms, which can be absorbed into the energy estimates. And applying the exterior derivative to (69) yields exactly (79) with a solution reconstructed also from
@@
It is inferred that the microlocal lift of the Navier-Stokes equations produces a well-defined dissipative dynamical system on the cosphere bundle, whose solution admits an explicit representation through evolution operators. The exact compatibility between the velocity and vorticity equations guaranties the consistency of the formalism.
This approach reduces the global regularity problem to the control and convergence of geometric microlocal invariants, the analysis of which is addressed in the following section.
@@ V Stability, Energy and Geometric Criteria for Regularity and Smoothness
For those reasons, throughout this section, $(M,g)$ denotes a smooth Riemannian manifold and $S^{*}M$ its unit cosphere bundle endowed with the canonical Liouville measure $\mu_{S^{*}M}$. We assume that the velocity field satisfies
$u\in L^{\infty}([t_{0},t];W^{2,\infty}(M))$ (88)
and that all geometric quantities associated with the effective connection $\nabla^{u}$ remain smooth on $t_{0},t]\times S^{*}M$. Let $\lambda\in\{\phi,\omega\}$ denote either the velocity $1$-form or the vorticity $2$-form on $M$, and $\tilde{\lambda}$ its microlocal lift to $S^{*}M$. All contractions and norms are taken with respect to the effective metric $g^{\prime}$.
Defining the fundamental local geometric quantity governing the analysis, the microlocal amplitude of $\tilde{\lambda}$ is defined pointwise by the contraction
$\|\tilde{\lambda}\|_{g^{\prime}}^{2}:=\langle\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}=\iota_{\tilde{\lambda}^{\sharp}}\tilde{\lambda}$ (89)
where $\tilde{\lambda}^{\sharp}$ denotes the metric dual with respect to $g^{\prime}$. This quantity measures the intensity of the lifted differential form along a given cotangent direction. Unlike classical norms, it retains full directional information encoded in the cosphere bundle. The corresponding global geometric functional is obtained by pairing this density with the canonical Liouville volume form on $S^{*}M$,
@@
where $\alpha$ is the contact form induced from the Liouville $1$-form on $T^{*}M$.
V.1 Energy Functionals
The microlocal evolution of $\tilde{\lambda}$ is governed by the equation
$\partial_{t}\tilde{\lambda}+\mathcal{L}_{X_{u}}\tilde{\lambda}+\nu\|\xi\|_{g^{\prime}}^{2}\,\tilde{\lambda}=\tilde{R}[\tilde{\lambda}]+\tilde{C}[\tilde{\lambda}]+\tilde{F}$ (91)
where $X_{u}$ denotes the canonical lifted transport vector field on $S^{*}M$, $\tilde{R}$ is a zeroth-order curvature operator induced by the effective connection, $\tilde{C}$ collects commutator terms arising from microlocalization, and $\tilde{F}$ represents external forcing. Doing a energy functional approach to this equation to have boundary conditions, we define the lifted quadratic energy functional as
$\mathcal{E}(t):=\frac{1}{2}\int_{S^{*}M}\langle\tilde{\lambda}(t),\tilde{\lambda}(t)\rangle_{g^{\prime}}\,\mu_{S^{*}M}$ (92)
Differentiating in time along solutions of (91) yields
$\frac{d}{dt}\mathcal{E}(t)$ $=\int_{S^{*}M}\langle\partial_{t}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}+\frac{1}{2}\int_{S^{*}M}\partial_{t}\langle\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}$ (93)
If the effective metric $g^{\prime}$ is fixed or depends smoothly on $u$, the second term is controlled by $\|\nabla u\|_{L^{\infty}}$ and can be absorbed into the estimates below.
Using (91), we estimate each contribution separately. The transport term satisfies
$\left\|\int_{S^{*}M}\langle\mathcal{L}_{X_{u}}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}\right\|\leq\sigma\|\nabla u\|_{L^{\infty}}\|\tilde{\lambda}\|_{L^{2}(S^{*}M)}^{2}$ (94)
while the dissipative term is nonnegative,
$\int_{S^{*}M}\langle\nu\|\xi\|_{g^{\prime}}^{2}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}=\nu\int_{S^{*}M}\|\xi\|_{g^{\prime}}^{2}\|\tilde{\lambda}\|_{g^{\prime}}^{2}\,\mu_{S^{*}M}\geq 0$ (95)
The curvature term admits the pointwise bound
$\|\langle\tilde{R}[\tilde{\lambda}],\tilde{\lambda}\rangle_{g^{\prime}}\|\leq\|\operatorname{Ric}^{u}\|_{L^{\infty}(S^{*}M)}\|\tilde{\lambda}\|_{g^{\prime}}^{2}$ (96)